OpenStudy (anonymous):

write an expression for f(x+h)-f(x) / h and then evaluate the result by putting h=0 1. f(x) = 3x 2. f(x) = -3x 3. f(x) = 3x + 10 4. f(x) = x^2 5. f(x) = x^2 + 15 6 f(x) = -x^2 7. f(x) = -x^2 + 30 8.f(x) = x^ + 5x 9. f(x) = x^2 + 5x + 10 10. f(x) = 8

OpenStudy (anonymous):

that's a lot....

OpenStudy (anonymous):

yup Im stump as to which direction to take

OpenStudy (anonymous):

the answer is the straight-up derivative. if you just want the answer without knowing how it's simple enough

OpenStudy (anonymous):

agree!

OpenStudy (anonymous):

@ML350 do you know derivatives or ur still in limits?

OpenStudy (anonymous):

all the problems can be done using these steps: 1. find f(x+h) 2. subtract f(x) from #1. 3. divide #1 and #2 by h 4. simplify. here's a sample using the first problem f(x) = 3x. apply step #1 to get \[f(x+h) = 3 (x+h) = 3x + 3h\] now subtract f(x) from it to get \[f(x+h)-f(x) = 3x +3h - 3x = 3h\] now divide both sides by h to get \[[f(x+h)-f(x)]/h = 3h/h = 3\] and that is the solution.

OpenStudy (anonymous):

heres another sample using problem #9: \[f(x+h) = (x+h)^2 + 5(x+h) + 10 = x^2 + 2hx + h^2 +5x + 5h + 10\] now simplify to get \[f(x+h) = x^2 + 5x + 2hx + 5h + h^2 + 10\] now subtract f(x) from it to get \[f(x+h)-f(x) = x^2 +5x + 2hx + 5h + h^2 + 10 -(x^2 + 5x + 10)\] simplifying, we get the following \[f(x+h) - f(x) = 2hx + 5h \] now divide both sides by an h to get \[[f(x+h)-f(x)]/h = (2hx + 5h)/h = 2x + 5\] which is the solution to the problem.

OpenStudy (anonymous):

let me know if you need further assistance

OpenStudy (anonymous):

Thank YOU! gonna get to these questions now....